7,9,-1,5, n what is n?
Answer by Edwin McCravy(20056)

7,9,-1,5, n what is n?

One of many possible general equations for a sequence that has
those first 4 terms is:

ax = (14x3-102x2+214x-105)/3  

Substitute x=1

a1 = [14(1)3-102(1)2+214(1)-105]/3 =

[14(1)-102(1)+214(1)-105]/3 = [14-102+214-105]/3 = 

21/3 = 7

Substitute x=2

a2 = [14(2)3-102(2)2+214(2)-105]/3 =

[14(8)-102(4)+214(2)-105]/3 = [112-408+428-105]/3 = 

27/3 = 9

Substitute x=3

a3 = [14(3)3-102(3)2+214(3)-105]/3 =

[14(27)-102(9)+214(3)-105]/3 = [378-918+642-105]/3 = 

-3/3 = -1

Substitute x=4

a4 = [14(4)3-102(4)2+214(4)-105]/3 =

[14(64)-102(16)+214(4)-105]/3 = [896-1632+856-105]/3 = 

15/3 = 5

Substitute x=5

a5 = [14(5)3-102(5)2+214(5)-105]/3 =

[14(125)-102(25)+214(5)-105]/3 = [1750-2550+1070-105]/3 = 

165/3 = 55

Answer: n = 55.

Edwin